OpenStudy (anonymous):

Solve the following system of equations. 3x – 2y + z = –6 x – 3y + z = –3 2x + y – 2z = –7 (1, 5, 1) (–2, 1, 2) (–3, –6, –7) (1, –3, 3)

OpenStudy (dumbcow):

do you remember how to solve a 2by2 system using elimination?

OpenStudy (dumbcow):

or addition method

OpenStudy (dumbcow):

the short easy way is by using matrix algebra and a calculator...but i doubt you have learned that yet

OpenStudy (anonymous):

no i haven't learned that

OpenStudy (dumbcow):

ok there are multiple steps... first you must eliminate a variable, meaning the goal is to reduce this to a system of 2 equations lets eliminate z, so pick any 2 equations and use elimination method to make the z's cancel out 3x -2y +z = -6 x -3y +z = -3 multiply bottom equation by -1 so that the "z" coefficients will add to zero then add the equations 3x -2y +z = -6 -x +3y -z = 3 -------------- 2x +y = -3 Now pick 2 other equations and do the same thing, eliminate z x -3y +z = -3 2x + y -2z = -7 multiply top equation by 2 so "z" coefficients add to zero 2x -6y +2z = -6 2x +y -2z = -7 -------------- 4x -5y = -13

OpenStudy (dumbcow):

Next step, take those 2 new equations and try to eliminate either x or y lets do y 2x +y = -3 4x -5y = -13 multiply top by 5 so that "y" coefficients add to zero 10x +5y = -15 4x -5y = -13 --------------- 14x = -28 Leaving a simple single equation to solve for x

OpenStudy (anonymous):

Oh ok and then you would plug x in to get y and z. Thank you so much for the help!

OpenStudy (dumbcow):

no problem